Secret:

Perform the indicated operation. (9z^3/16xy) * (4x/27z^3)

dude:

\((9\cdot \frac{z^3}{16xy})(4\cdot \frac{x}{27z^3})\)?

kittybasil:

Assuming the previously mentioned expression: \[\large{(9\cdot\frac{z^3}{16xy})(4\cdot\frac{x}{27z^3})}\] Isolate the numerical values. \[\large{(\frac{9}{16}\cdot\frac{z^3}{xy})(\frac{4}{27}\cdot\frac{x}{z^3})}\] Find similar terms in numerator and denominator that you can cancel. I crossed them out so you can see. \[\large{(\frac{9}{16}\cdot\frac{\cancel{z^3}}{\bcancel{x}y})(\frac{4}{27}\cdot\frac{\bcancel{x}}{\cancel{z^3}})}\] So now you have this left: \[\large{(\frac{9}{16}\cdot\frac{1}{y})(\frac{4}{27})}\] Simplify anything you can with the numbers. I rearranged them to show relationships. \[\large{\frac{9}{27}\cdot\frac{4}{16}\cdot\frac{1}{y}}\rightarrow\frac{1}{3}\cdot\frac{1}{4}\cdot\frac{1}{y}=\frac{1}{12y}\] And that's your answer. Sorry if it's weird, I haven't been on here in awhile so I forgot how to use the bracket-type LaTeX formatting.